It is integration problem.... Solve problem.........

Grade:12

4 Answers

L GOUSE BASHA
7 years ago
P { margin-bottom: 0.08in; }H2 { margin-bottom: 0.08in; }H2.western { font-family: "Liberation Sans",sans-serif; font-size: 14pt; font-style: italic; }H2.cjk { font-family: "WenQuanYi Zen Hei"; font-size: 14pt; font-style: italic; }H2.ctl { font-family: "Lohit Devanagari"; font-size: 14pt; font-style: italic; }

X2/4-X/4+7/8ln|2X+3|+C

RAKESH CHINDAM
7 years ago
P { margin-bottom: 0.08in; }H2 { margin-bottom: 0.08in; }H2.western { font-family: "Liberation Sans",sans-serif; font-size: 14pt; font-style: italic; }H2.cjk { font-family: "WenQuanYi Zen Hei"; font-size: 14pt; font-style: italic;

X2/4-X/4+7/8ln|2X+3|+C

SHANMUKESHWAR
7 years ago
x^2+x+1/2x+3=(x/2-1/4)+7/4*(2x+3)=x^2/4-x/4+7/4*ln(2x+3)/2+c where c is a arbitary constant
 
x^2/4-x/4+7/8*ln(2x+3)/2+c.
T.kumar
7 years ago
my answer is:
 
                       x^2/4-x/4+7/8*ln(2x+3)/2+c.
all the best for your bright future..............this is my answeer..